OpenStudy (anonymous):

Find an equation for the line with the given properties. Express the answer using slope intercept form of the equation of a line. Slope = (-9/2); containing the point (8, -5).

OpenStudy (aroub):

y=mx+b , m= -9/2 , (8,-5) -5=-9/2(8) +b -5=-72/2 +b 72/2-10/2 =b 62/2=b 31=b so the equation is y= -9/2 +31

OpenStudy (anonymous):

Thanks :)

OpenStudy (aroub):

anytime!! :D